OpenStudy (anonymous):

Solve the following equation: (5x - 4)(9x + 7) = 0

OpenStudy (anonymous):

solve for x make one binomial=0, like separate them for example. (5x-4)=0, and than (9x+7)=0.5x-4=0=5x=4, than divide that 5 to get it away from the x so than its x=4/5, and you do the same way for 9x+7=0, get 7 to the other side so, 9x=-7 than divide 9 to remove it from x so than the answer is x=-7/9

OpenStudy (anonymous):